OpenStudy (anonymous):

can anyone explain this to me in a simple way please. if a third term of an arithmetic sequence is 14 and the 8th term is -1, find the values of a1 and d of the sequence

OpenStudy (anonymous):

hey dude didnt u got it?????

OpenStudy (anonymous):

not yet I will write my solution then you check where I got it wrong.

OpenStudy (anonymous):

ok go on

OpenStudy (anonymous):

i gave u the correct explanation in the last thread bro once check it and then do it again

OpenStudy (anonymous):

a3=14, a8=-1 a3=a1+(n-1)d 14=a1+(3-1)d a1=14-2d equation 1

OpenStudy (anonymous):

yeah perfect

OpenStudy (anonymous):

ok now equation 2 a8=a1+(n-1)d -1=a1+(n-1)d a1=-1+7d equation 2

OpenStudy (anonymous):

and the second will a1=-1-7d

OpenStudy (anonymous):

Solving those two u will get a1=20 and d=-3

OpenStudy (anonymous):

just a sec why is it -7 not 7 cause 8-1=7

OpenStudy (anonymous):

arey yar +7 will be there when it is on left hand side and it will be negative when u transfer it to the other side

OpenStudy (anonymous):

actual transfer in sesnse is that u will subtract it on both sides on the left hand side u will remain with zero and on rhs u will get that -ve sign

OpenStudy (anonymous):

substitute same as u did in the first eqn a1+(8-1)d=-1 a1+7d=-1 a1=-1-7d ok????

OpenStudy (anonymous):

humm thanks that was my mistake I guess

OpenStudy (anonymous):

hmmm k k

OpenStudy (anonymous):

got d=-3

OpenStudy (anonymous):

a1=8 am I correct

OpenStudy (anonymous):

look at this I am trying to get the sequence but it does not work

OpenStudy (anonymous):

a1 is 20 dude

OpenStudy (anonymous):

yeah you are right sorry

OpenStudy (anonymous):

sorry forgot to give you a medal